the sides of a rectangle are in the ratio 3:2. what is the length of each side if the perimeter of te rectangle is 55 centimeters
Answer by lwsshak3(11628)

The sides of a rectangle are in the ratio 3:2. what is the length of each side if the perimeter of te rectangle is 55 centimeters
**
let x=length
let y=width
2x+2y=55
x/y=3/2
3y=2x
y=2x/3
..
2x+2(2x/3)=55
2x+4x/3=55
6x+4x=165
10x=165
x=16.5
y=2*16.5/3=11
..
Ans:
length=2x=33
width=2y=22